# APCTP-Pre2023-010

### v2

@@ IV Criteria for the growth of perturbations
Having explored the two-field case we move to investigate the growth of perturbations for an arbitrary number of fields. The set of the $2\mathcal{N}$ Hamilton’s equations can be written in the matrix form $X^{\prime}=JX$ with $X\equiv(\mathcal{R},q_{i},P_{\mathcal{R}},P_{i})^{T}$ and $J$ being a $2\mathcal{N}\times 2\mathcal{N}$ time-dependent matrix. The behaviour of perturbations is qualitatively different on sub/super-horizon scales so we will study the two regimes separately.
IV.1 Subhorizon evolution
@@
$\mathcal{R}_{,x}$ $=p_{\mathcal{R}}-\sqrt{\frac{2}{\epsilon}}{\Omega\over x}q_{\rm n}\,,$ (105)
$p_{\mathcal{R},x}$ $={2\over x}p_{\mathcal{R}}-\mathcal{R}\,,$ (106)
$q_{i,x}$ $=p_{i}-{Z_{ij}\over x}q^{j}\,,$ (107)
$p_{i,x}$ $={2\over x}p_{i}+{2\sqrt{2\epsilon}\Omega\over x}p_{\mathcal{R}}\delta_{i{\rm n}}-{Z_{ij}\over x}p^{j}-q_{i}-{\mathcal{M}_{ij}\over x^{2}}q^{j}-{3\Omega^{2}\over x^{2}}\delta_{i{\rm n}}q_{\rm n}\,,$ (108)
where the indices $i$ and $j$ stand for the $\mathcal{N}-1$ isocurvature perturbations orthogonal to $\mathcal{R}$, but $a$ and $b$ denote all the $\mathcal{N}$ perturbations, and $P_{a}=-x\,p_{a}$. Moreover, we have ignored slow-roll corrections, i.e. $\epsilon,|\eta|\ll 1$. Similar to section III.2 we split $J$ in two parts, $J=J_{(0)}+J_{(1)}$. Considering large bending parameters $\Omega,T,\cdots$ only the friction terms $2p_{a}/x$ fall in $J_{(1)}$ and thus we neglect the friction terms and focus on $J_{(0)}$. Next, we introduce two enlarged, $\mathcal{N}\times\mathcal{N}$ matrices ~$\tilde{\mathcal{M}}_{ab}$ and $\tilde{Z}_{ab}$ defined as
~$\tilde{\mathcal{M}}_{ab}\equiv\begin{pmatrix}0&0\\ 0&\mathcal{M}_{ij}+3\Omega^{2}\delta_{i\rm n}\delta_{j\rm n}\\ \end{pmatrix}\quad\text{and}\quad\tilde{Z}_{ab}\equiv\begin{pmatrix}0&0\\ 0&Z_{ij}\\ \end{pmatrix}\,,$ (109)
@@
$J_{(0)}=\begin{pmatrix}-\dfrac{1}{x}\tilde{Z}_{ab}-\dfrac{1}{x}\sqrt{\dfrac{2}{\epsilon}}\Omega\delta_{a\mathcal{R}}\delta_{b\rm n}&I_{ab}\vspace{0.3em}\\ -I_{ab}-\dfrac{1}{x^{2}}\tilde{\mathcal{M}}_{ab}&-\dfrac{1}{x}\tilde{Z}_{ab}+\dfrac{2}{x}\sqrt{2\epsilon}\Omega\delta_{a\rm n}\delta_{b\mathcal{R}}\end{pmatrix}\equiv\begin{pmatrix}A&B\\ C&D\end{pmatrix}\,.$ (110)
@@
$\lambda^{2\mathcal{N}}-\text{Tr}(J_{(0)})\lambda^{2\mathcal{N}-1}+\dots+\det(J_{(0)})=0\,,$ (111)
@@
$\text{det}(J_{(0)})=\text{det}(A)\det\left(D-CA^{-1}B\right)=\det\left(DA-C\right)\,,$ (112)
@@
$DA-C=I_{ab}+{1\over x^{2}}\left(\tilde{\mathcal{M}}_{ab}-4\Omega^{2}\delta_{a\rm n}\delta_{b\rm n}+\tilde{Z}_{ak}\tilde{Z}_{kb}\right)\,.$ (113)
@@
$\det(J_{(0)})=\det\left[I_{ij}+{1\over x^{2}}\left(\mathcal{M}_{ij}-4\Omega^{2}\delta_{i\rm n}\delta_{j\rm n}+\tilde{Z}_{ik}Z_{kj}\right)\right]\equiv\det\left(I+{1\over x^{2}}\mathcal{M}_{\rm sub}\right)\,,$ (114)
where $\mathcal{M}_{\rm sub}$ generalizes the effective mass of the isocurvature perturbation on subhorizon scales, and the indices $i$ and $j$ only count for these $\mathcal{N}-1$ isocurvature perturbations. With the expression of the determinant to our disposal we are ready to predict when the perturbations will be amplified. Using the eigenvalues $\mu_{i}$ of the matrix $\mathcal{M}_{\rm sub}$, we can factorize the determinant in the following way:
$\det(J_{(0)})=\prod_{i}\left(1+{\mu_{i}\over x^{2}}\right)\,.$ (115)
@@
$\det(J_{(0)})=\prod_{m}\left(1+{\mu_{m}\over x^{2}}\right)\prod_{j}\left(1+{2\over x^{2}}\Re(\mu_{j})+{|\mu_{j}|^{2}\over x^{4}}\right)\,,$ (116)
@@
We also mention that the eigenvalue equation (111) can be obtained from the determinant of the following matrix:
$\det(J_{\rm eig,sub})\equiv\det\left(\lambda^{2}I-(A+D)\lambda+DA-C\right)=0\,.$ (117)
@@
$\det(J_{\rm eig,sub})=$ $\left(\lambda^{2}+1\right)\det\left[(\lambda^{2}+1)I+{2\over x}Z\lambda+{1\over x^{2}}\mathcal{M}_{\rm sub}\right]+{4\Omega^{2}\lambda^{2}\over x^{2}}\det\left(\left[(\lambda^{2}+1)I+{2\over x}Z\lambda+{1\over x^{2}}\mathcal{M}_{\rm sub}\right]_{\rm nn}\right)=0\,.$ (118)
In the case of two fields this equation reproduces the four eigenvalues mentioned in (98) and (99) of section III.2 after setting the cofactor $[\cdots]_{\rm nn}$ to 1.
IV.2 Superhorizon evolution
On superhorizon scales the system of (48)-(51) is already in perturbative form and so we can isolate the $\kappa$-dependent part as
$X^{\prime}=J_{\rm sup(0)}X+J_{\rm sup(1)}X\,.$ (119)
@@ IV.2 Superhorizon evolution
$Y^{\prime}=J_{\rm eig(0)}Y+SJ_{\rm sup(1)}S^{-1}Y\,.$ (120)
The second term is proportional to $\kappa^{2}$ and remains finite when integrated in the interval $(0,\kappa_{i})$ and so (by Levinson’s theorem levinson ) stability can be inferred by focusing on the eigenvalues of the $\kappa$-independent part. Thus, we can set $\kappa=0$ in the original system and focus on the eigenvalues of $J_{\rm sup(0)}$. Due to the structure of $J_{\rm sup(0)}$ the evolution equations for $\mathcal{R}$ and $P_{\mathcal{R}}$ decouple from the rest equations so we can focus entirely on the orthogonal subspace. The stability matrix describing the orthogonal perturbations is explicitly
$J_{\rm orth,sup}=\begin{pmatrix}Z_{ij}&I_{ij}\vspace{0.3em}\\ -\mathcal{M}_{ij}-3\Omega^{2}\delta_{i\rm n}\delta_{j\rm n}&-(3-\epsilon)I_{ij}+Z_{ij}\end{pmatrix}\,,$ (121)
@@
$\det(J_{\rm eig,sup})=\det\left(\lambda^{2}I+\left((3-\epsilon)I-2Z\right)\lambda+\mathcal{M}_{\rm sup}\right)=0\,,$ (122)
@@
$\left(\mathcal{M}_{\rm sup}\right)_{ij}\equiv\mathcal{M}_{ij}+3\Omega^{2}\delta_{i\rm n}\delta_{j\rm n}+Z_{ik}Z_{kj}-(3-\epsilon)Z_{ij}\,.$ (123)
@@
$\left[\lambda^{2}+(3-\epsilon)\lambda+\mathcal{M}_{\rm nn}+3\Omega^{2}-T^{2}\right]\det\left(\left[J_{\rm eig,sup}\right]_{\rm nn}\right)-\left[\mathcal{M}_{\rm nb}^{2}-\left(2\lambda+3-\epsilon\right)^{2}T^{2}\right]\det\left(\left[J_{\rm eig,sup}\right]_{\rm nn,bb}\right)=0\,.$ (124)
To achieve stability on superhorizon scales the eigenvalues of the this equation should have negative real part.
@@
$\lambda_{1,2}=-{1\over 2}\left[3-\epsilon\pm\sqrt{(3-\epsilon)^{2}-4m_{\rm eff}}\right]\,,$ (125)
@@
$\lambda_{i}=-{1\over 2}\left\{3-\epsilon\pm\sqrt{(3-\epsilon)^{2}-2\text{Tr}(\mathcal{M}_{\rm sup})-8T^{2}\pm 2\sqrt{\left[\text{Tr}(\mathcal{M}_{\rm sup})+4T^{2}\right]^{2}-4\det(\mathcal{M}_{\rm sup})}}\right\}\,.$ (126)
Demanding $\Re(\lambda_{i})<0$ we find the conditions for stable perturbations listed in Christodoulidis:2022vww . Similar to the turn rate, the presence of the torsion provides a stabilizing effect on the orthogonal perturbations.
V Models with isometries
Models with isometries have attracted a lot of attention in the literature. Not only they yield simpler models, but they are appropriate candidates of rapid-turn or rapid-twist behaviour Christodoulidis:2022vww . Moreover, the isometry structure linearly relates the nb-component of the mass matrix with the torsion through $\mathcal{M}_{\rm nb}=(3-\epsilon)T$. In this section we will demonstrate that the bb-component satisfies $\mathcal{M}_{\rm bb}=T^{2}$, which implies that for three fields the mass matrix has only one free parameter that does not automatically depend on the kinematical quantities, and then we will focus on the case of massless isocuvature perturbations.
V.1 $\mathcal{N}-1$ isometries
@@
$Tb^{i}\approx\Gamma^{i}_{\chi j}v^{j}+\Omega t^{i}\,,$ (127)
@@
$T^{2}+\Omega^{2}=\Gamma^{i}_{\chi k}\Gamma^{j}_{\chi l}v^{k}v^{l}\,.$ (128)
@@
${\,\mathrm{d}\phi^{i}\over\,\mathrm{d}\phi^{j}}={V^{,i}\over V^{,j}}\,,$ (130)
@@
$M_{\rm bb}=V_{\rm bb}-R_{kijl}\dot{\phi}^{i}\dot{\phi}^{j}b^{k}b^{l}\approx-b^{k}b^{l}\Gamma_{kl}^{\chi}V_{,\chi}-G_{kd}\left(\Gamma^{d}_{i\chi}\Gamma^{\chi}_{lj}-\Gamma^{d}_{l\chi}\Gamma^{\chi}_{ij}\right)\dot{\phi}^{i}\dot{\phi}^{j}b^{k}b^{l}\approx T^{2}H^{2}\,,$ (131)
where we made use of (127) and the vanishing of the effective gradient $V_{,\chi}+\Gamma^{\chi}_{ij}\dot{\phi}^{i}\dot{\phi}^{j}\approx 0$.
Because of this form of the components of the mass matrix, the $x$-dependent eigenvalues which follow from (118), greatly simplify up to slow-roll corrections. For three fields they are given by
$\lambda_{\pm\pm}$ =±$=\pm{1\over\sqrt{2}x}\sqrt{-2x^{2}-\mu_{\rm eff}\pm\sqrt{\mu_{\rm eff}^{2}+16x^{2}\left(T^{2}+\Omega^{2}\right)+4(3-\epsilon)^{2}T^{2}}}\,,$ (132)
$\lambda_{5,6}$ =±$=\pm i\,,$ (133)
@@
$\mu_{\rm eff}\equiv\mathcal{M}_{\rm nn}+3\Omega^{2}+3T^{2}\,.$ (134)
@@
$x^{2}+\mathcal{M}_{\rm nn}-\Omega^{2}-T^{2}-{(3-\epsilon)^{2}T^{2}\over x^{2}}<0\,,$ (135)
which, for relatively large values of $\Omega$ and $T$, marks the onset of the amplification (see figure 3). Notice that the last term in (135) is suppressed on subhorizon scales and, hence, this three-field model is very similar to a two-field model with a turn rate given by $\Omega^{2}_{\rm 2f}=\Omega^{2}+T^{2}$; this was also noticed in Aragam:2023adu for similar choices of the mass matrix elements but with $\mathcal{M}_{\rm nb}=0$, which is the reason why the last term in (132) was absent. On superhorizon scales by investigating (126), the condition for stability reduces to simply $\mu_{\rm eff}>0$.
V.2 A toy model with massless isocurvature perturbations
As we saw in section IV.2 on superhorizon scales the orthogonal perturbations admit exponentially decaying solutions with exponents given by (126). We will call the perturbations massless when all eigenvalues become zero. For three fields, in particular, the two eigenvalues become zero only if
$\det(\mathcal{M}_{\rm sup})=0\quad\text{and}\quad\text{Tr}(\mathcal{M}_{\rm sup})+4T^{2}=0\,.$ (136)

### v3

@@ 4 Criteria for the growth of perturbations
Having explored the two-field case we move on to investigate the growth of perturbations for an arbitrary number of fields. The set of the $2\mathcal{N}$ Hamilton’s equations can be written in the matrix form $X^{\prime}=JX$ with $X\equiv(\mathcal{R},q_{i},P_{\mathcal{R}},P_{i})^{T}$ and $J$ being a $2\mathcal{N}\times 2\mathcal{N}$ time-dependent matrix. The behaviour of perturbations is qualitatively different on sub/super-horizon scales so we will study the two regimes separately.
4.1 Subhorizon evolution
@@
$\mathcal{R}_{,x}$ $=p_{\mathcal{R}}-\sqrt{\frac{2}{\epsilon}}{\Omega\over x}q_{\rm n}\,,$ (4.1)
$p_{\mathcal{R},x}$ $={2\over x}p_{\mathcal{R}}-\mathcal{R}\,,$ (4.2)
$q_{i,x}$ $=p_{i}-{Z_{ij}\over x}q^{j}\,,$ (4.3)
$p_{i,x}$ $={2\over x}p_{i}+{2\sqrt{2\epsilon}\Omega\over x}p_{\mathcal{R}}\delta_{i{\rm n}}-{Z_{ij}\over x}p^{j}-q_{i}-{\mathcal{M}_{ij}\over x^{2}}q^{j}-{3\Omega^{2}\over x^{2}}\delta_{i{\rm n}}q_{\rm n}\,,$ (4.4)
where the indices $i$ and $j$ stand for the $\mathcal{N}-1$ isocurvature perturbations orthogonal to $\mathcal{R}$, but $a$ and $b$ denote all the $\mathcal{N}$ perturbations, and $P_{a}=-x\,p_{a}$. Moreover, we have ignored slow-roll corrections, i.e. $\epsilon,|\eta|\ll 1$. Similar to section 3.2 we split $J$ in two parts, $J=J_{(0)}+J_{(1)}$. Considering large bending parameters $\Omega,T,\cdots$ only the friction terms $2p_{a}/x$ fall in $J_{(1)}$ and thus we neglect the friction terms and focus on $J_{(0)}$. Next, we introduce two enlarged, $\mathcal{N}\times\mathcal{N}$ matrices ~$\tilde{\mathcal{M}}_{ab}$ and $\tilde{Z}_{ab}$ defined as
~$\tilde{\mathcal{M}}_{ab}\equiv\begin{pmatrix}0&0\\ 0&\mathcal{M}_{ij}+3\Omega^{2}\delta_{i\rm n}\delta_{j\rm n}\\ \end{pmatrix}\quad\text{and}\quad\tilde{Z}_{ab}\equiv\begin{pmatrix}0&0\\ 0&Z_{ij}\\ \end{pmatrix}\,,$ (4.5)
@@
$J_{(0)}=\begin{pmatrix}-\dfrac{1}{x}\tilde{Z}_{ab}-\dfrac{1}{x}\sqrt{\dfrac{2}{\epsilon}}\Omega\delta_{a\mathcal{R}}\delta_{b\rm n}&I_{ab}\vspace{0.3em}\\ -I_{ab}-\dfrac{1}{x^{2}}\tilde{\mathcal{M}}_{ab}&-\dfrac{1}{x}\tilde{Z}_{ab}+\dfrac{2}{x}\sqrt{2\epsilon}\Omega\delta_{a\rm n}\delta_{b\mathcal{R}}\end{pmatrix}\equiv\begin{pmatrix}A&B\\ C&D\end{pmatrix}\,.$ (4.6)
@@
$\lambda^{2\mathcal{N}}-\text{Tr}(J_{(0)})\lambda^{2\mathcal{N}-1}+\dots+\det(J_{(0)})=0\,,$ (4.7)
@@
$\text{det}(J_{(0)})=\text{det}(A)\det\left(D-CA^{-1}B\right)=\det\left(DA-C\right)\,,$ (4.8)
@@
$DA-C=I_{ab}+{1\over x^{2}}\left(\tilde{\mathcal{M}}_{ab}-4\Omega^{2}\delta_{a\rm n}\delta_{b\rm n}+\tilde{Z}_{ak}\tilde{Z}_{kb}\right)\,.$ (4.9)
@@
$\det(J_{(0)})=\det\left[I_{ij}+{1\over x^{2}}\left(\mathcal{M}_{ij}-4\Omega^{2}\delta_{i\rm n}\delta_{j\rm n}+\tilde{Z}_{ik}Z_{kj}\right)\right]\equiv\det\left(I+{1\over x^{2}}\mathcal{M}_{\rm sub}\right)\,,$ (4.10)
where $\mathcal{M}_{\rm sub}$ generalizes the effective mass of the isocurvature perturbation on subhorizon scales, and the indices $i$ and $j$ only count for these $\mathcal{N}-1$ isocurvature perturbations. With the expression of the determinant at our disposal we are ready to predict when the perturbations will be amplified. Using the eigenvalues $\mu_{i}$ of the matrix $\mathcal{M}_{\rm sub}$, we can factorize the determinant in the following way:
$\det(J_{(0)})=\prod_{i}\left(1+{\mu_{i}\over x^{2}}\right)\,.$ (4.11)
@@
$\det(J_{(0)})=\prod_{m}\left(1+{\mu_{m}\over x^{2}}\right)\prod_{j}\left(1+{2\over x^{2}}\Re(\mu_{j})+{|\mu_{j}|^{2}\over x^{4}}\right)\,,$ (4.12)
@@
We also mention that the eigenvalue equation (4.7) can be obtained from the determinant of the following matrix:
$\det(J_{\rm eig,sub})\equiv\det\left(\lambda^{2}I-(A+D)\lambda+DA-C\right)=0\,.$ (4.13)
@@
$\det(J_{\rm eig,sub})=$ $\left(\lambda^{2}+1\right)\det\left[(\lambda^{2}+1)I+{2\over x}Z\lambda+{1\over x^{2}}\mathcal{M}_{\rm sub}\right]$ (4.14)
$+{4\Omega^{2}\lambda^{2}\over x^{2}}\det\left(\left[(\lambda^{2}+1)I+{2\over x}Z\lambda+{1\over x^{2}}\mathcal{M}_{\rm sub}\right]_{\rm nn}\right)=0\,.$
In the case of two fields this equation reproduces the four eigenvalues mentioned in (3.47) and (3.48) of section 3.2 after setting the cofactor $[\cdots]_{\rm nn}$ to 1.
4.2 Superhorizon evolution
On superhorizon scales the system of (2.48)-(2.51) is already in perturbative form and so we can isolate the $\kappa$-dependent part as
$X^{\prime}=J_{\rm sup(0)}X+J_{\rm sup(1)}X\,.$ (4.15)
@@ 4.2 Superhorizon evolution
$Y^{\prime}=J_{\rm eig(0)}Y+SJ_{\rm sup(1)}S^{-1}Y\,.$ (4.16)
The second term is proportional to $\kappa^{2}$ and remains finite when integrated in the interval $(0,\kappa_{i})$ and so (by Levinson’s theorem [71]) stability can be inferred by focusing on the eigenvalues of the $\kappa$-independent part. Thus, we can set $\kappa=0$ in the original system and focus on the eigenvalues of $J_{\rm sup(0)}$. Due to the structure of $J_{\rm sup(0)}$ the evolution equations for $\mathcal{R}$ and $P_{\mathcal{R}}$ decouple from the rest equations so we can focus entirely on the orthogonal subspace. The stability matrix describing the orthogonal perturbations is explicitly
$J_{\rm orth,sup}=\begin{pmatrix}Z_{ij}&I_{ij}\vspace{0.3em}\\ -\mathcal{M}_{ij}-3\Omega^{2}\delta_{i\rm n}\delta_{j\rm n}&-(3-\epsilon)I_{ij}+Z_{ij}\end{pmatrix}\,,$ (4.17)
@@
$\det(J_{\rm eig,sup})=\det\left(\lambda^{2}I+\left((3-\epsilon)I-2Z\right)\lambda+\mathcal{M}_{\rm sup}\right)=0\,,$ (4.18)
@@
$\left(\mathcal{M}_{\rm sup}\right)_{ij}\equiv\mathcal{M}_{ij}+3\Omega^{2}\delta_{i\rm n}\delta_{j\rm n}+Z_{ik}Z_{kj}-(3-\epsilon)Z_{ij}\,.$ (4.19)
@@
[…]
To achieve stability on superhorizon scales the eigenvalues of this equation should have negative real part.
@@
$\lambda_{1,2}=-{1\over 2}\left[3-\epsilon\pm\sqrt{(3-\epsilon)^{2}-4m_{\rm eff}}\right]\,,$ (4.21)
@@
$\lambda_{i}=-{1\over 2}\left\{3-\epsilon\pm\sqrt{(3-\epsilon)^{2}-2\text{Tr}(\mathcal{M}_{\rm sup})-8T^{2}\pm 2\sqrt{\left[\text{Tr}(\mathcal{M}_{\rm sup})+4T^{2}\right]^{2}-4\det(\mathcal{M}_{\rm sup})}}\right\}\,.$ (4.22)
Demanding $\Re(\lambda_{i})<0$ we find the conditions for stable perturbations listed in [72]. Similar to the turn rate, the presence of the torsion provides a stabilizing effect on the orthogonal perturbations.
5 Models with isometries
Models with isometries have attracted a lot of attention in the literature. Not only they yield simpler models, but they are appropriate candidates of rapid-turn or rapid-twist behaviour [72]. Moreover, the isometry structure linearly relates the nb-component of the mass matrix with the torsion through $\mathcal{M}_{\rm nb}=(3-\epsilon)T$. In this section we will demonstrate that the bb-component satisfies $\mathcal{M}_{\rm bb}=T^{2}$, which implies that for three fields the mass matrix has only one free parameter that does not automatically depend on the kinematical quantities, and then we will focus on the case of massless isocuvature perturbations.
5.1 $\mathcal{N}-1$ isometries
@@
$Tb^{i}\approx\Gamma^{i}_{\chi j}v^{j}+\Omega t^{i}\,,$ (5.1)
@@
$T^{2}+\Omega^{2}=\Gamma^{i}_{\chi k}\Gamma^{j}_{\chi l}v^{k}v^{l}\,.$ (5.2)
@@
${\,\mathrm{d}\phi^{i}\over\,\mathrm{d}\phi^{j}}={V^{,i}\over V^{,j}}\,,$ (5.4)
@@
$M_{\rm bb}=V_{\rm bb}-R_{kijl}\dot{\phi}^{i}\dot{\phi}^{j}b^{k}b^{l}\approx-b^{k}b^{l}\Gamma_{kl}^{\chi}V_{,\chi}-G_{kd}\left(\Gamma^{d}_{i\chi}\Gamma^{\chi}_{lj}-\Gamma^{d}_{l\chi}\Gamma^{\chi}_{ij}\right)\dot{\phi}^{i}\dot{\phi}^{j}b^{k}b^{l}\approx T^{2}H^{2}\,,$ (5.5)
where we made use of (5.1) and the vanishing of the effective gradient $V_{,\chi}+\Gamma^{\chi}_{ij}\dot{\phi}^{i}\dot{\phi}^{j}\approx 0$.
Because of this form of the components of the mass matrix, the $x$-dependent eigenvalues which follow from (4.14), greatly simplify up to slow-roll corrections. For three fields they are given by
$\lambda_{\pm\pm}$ =±$=\pm{1\over\sqrt{2}x}\sqrt{-2x^{2}-\mu_{\rm eff}\pm\sqrt{\mu_{\rm eff}^{2}+16x^{2}\left(T^{2}+\Omega^{2}\right)+4(3-\epsilon)^{2}T^{2}}}\,,$ (5.6)
$\lambda_{5,6}$ =±$=\pm i\,,$ (5.7)
@@
$\mu_{\rm eff}\equiv\mathcal{M}_{\rm nn}+3\Omega^{2}+3T^{2}\,.$ (5.8)
@@
$x^{2}+\mathcal{M}_{\rm nn}-\Omega^{2}-T^{2}-{(3-\epsilon)^{2}T^{2}\over x^{2}}<0\,,$ (5.9)
which, for relatively large values of $\Omega$ and $T$, marks the onset of the amplification (see figure 3). Notice that the last term in (5.9) is suppressed on subhorizon scales and, hence, this three-field model is very similar to a two-field model with a turn rate given by $\Omega^{2}_{\rm 2f}=\Omega^{2}+T^{2}$; this was also noticed in [45] for similar choices of the mass matrix elements but with $\mathcal{M}_{\rm nb}=0$, which is the reason why the last term in (5.6) was absent. On superhorizon scales by investigating (4.22), the condition for stability reduces to simply $\mu_{\rm eff}>0$.
5.2 A toy model with massless isocurvature perturbations
As we saw in section 4.2 on superhorizon scales the orthogonal perturbations admit exponentially decaying solutions with exponents given by (4.22). We will call the perturbations massless when all eigenvalues become zero. For three fields, in particular, the two eigenvalues become zero only if
$\det(\mathcal{M}_{\rm sup})=0\quad\text{and}\quad\text{Tr}(\mathcal{M}_{\rm sup})+4T^{2}=0\,.$ (5.10)